# RefLoRA: Refactored Low-Rank Adaptation for Efficient Fine-Tuning of Large Models

### v1

@@ B.1 Proof of Lemma 1
This suggests, there must exist an invertible matrix $\mathbf{P}_{t}\in^{r\times r}$ such that $\tilde{\mathbf{A}}_{t}=\mathbf{A}_{t}\mathbf{P}_{t}$.
@@
=$\overset{(a)}{=}-\eta\tilde{\mathbf{A}}_{t}\tilde{\mathbf{A}}_{t}^{\top}\nabla\ell(\mathbf{W}_{t})-\eta\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t}\tilde{\mathbf{B}}_{t}^{\top}+\eta^{2}\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t}\tilde{\mathbf{A}}_{t}^{\top}\nabla\ell(\mathbf{W}_{t})$
=$\overset{(b)}{=}-\eta\mathbf{A}_{t}\mathbf{P}_{t}\mathbf{P}_{t}^{\top}\mathbf{A}_{t}^{\top}\nabla\ell(\mathbf{W}_{t})-\eta\nabla\ell(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{P}_{t}^{-\top}\mathbf{P}_{t}^{-1}\mathbf{B}_{t}^{\top}+\eta^{2}\nabla\ell(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{A}_{t}^{\top}\nabla\ell(\mathbf{W}_{t})$
=$\overset{(c)}{=}\mathbf{A}_{t}(\mathbf{P}_{t}\mathbf{P}_{t}^{\top})\Delta\mathbf{B}_{t}^{\top}+\Delta\mathbf{A}_{t}(\mathbf{P}_{t}\mathbf{P}_{t}^{\top})^{-1}\mathbf{B}_{t}^{\top}+\Delta\mathbf{A}_{t}\Delta\mathbf{B}_{t}^{\top}$ (12)
where $(a)$ uses (3) and that $\mathbf{W}_{t}^{\mathrm{pre}}+\tilde{\mathbf{A}}_{t}\tilde{\mathbf{B}}_{t}^{\top}=\mathbf{W}_{t}^{\mathrm{pre}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}=\mathbf{W}_{t}$, $(b)$ is due to $\tilde{\mathbf{A}}_{t}=\mathbf{A}_{t}\mathbf{P}_{t}$ and $\tilde{\mathbf{B}}_{t}=\mathbf{B}_{t}\mathbf{P}^{-\top}$, and $(c)$ leverages (1).
@@ B.2 Proof of Proposition 2
=$\overset{(a)}{=}\ell(\mathbf{W}_{t})+\big{\langle}\nabla\ell(\mathbf{W}_{t}),\mathbf{A}_{t}\mathbf{S}_{t}\Delta\mathbf{B}_{t}^{\top}+\Delta\mathbf{A}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}-\Delta\mathbf{A}_{t}\Delta\mathbf{B}_{t}^{\top}\big{\rangle}_{\mathrm{F}}+$
@@
=$\overset{(b)}{=}\big{\langle}\nabla\ell(\mathbf{W}_{t}),\mathbf{A}_{t}\mathbf{S}_{t}\Delta\mathbf{B}_{t}^{\top}+\Delta\mathbf{A}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\big{\rangle}_{\mathrm{F}}+\frac{L}{2}\Big{[}\big{\|}\mathbf{A}_{t}\mathbf{S}_{t}\Delta\mathbf{B}_{t}^{\top}\big{\|}_{\mathrm{F}}^{2}+\big{\|}\Delta\mathbf{A}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\big{\|}_{\mathrm{F}}^{2}+$
@@
=$\overset{(c)}{=}\frac{L}{2}\Big{[}\Big{\|}\frac{1}{L}\nabla\ell(\mathbf{W}_{t})+\mathbf{A}_{t}\mathbf{S}_{t}\Delta\mathbf{B}_{t}^{\top}\Big{\|}_{\mathrm{F}}^{2}+\Big{\|}\frac{1}{L}\nabla\ell(\mathbf{W}_{t})+\Delta\mathbf{A}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\Big{\|}_{\mathrm{F}}^{2}+$ (13)
@@
When performing fine-tuning from a pretrained weight, both $\eta$ and $\|\nabla\ell(\mathbf{W}_{t})\|_{2}$ are observed to be tiny [^1] Typically, $\eta=\mathcal{O}(10^{-4})$, and $\|\nabla\ell(\mathbf{W}_{t})\|_{2}\leq\|\nabla\ell(\mathbf{W}_{t})\|_{\mathrm{F}}=\mathcal{O}(10^{-1})$ per matrix.. As a result, (B.2) is dominated by (B.2), and is neglectable in practice; see also [56, 63].
@@
=$\overset{(a)}{=}\frac{L\eta^{2}}{2}\|\nabla\ell(\mathbf{W}_{t})\|_{2}^{2}\Big{[}\big{\|}\mathbf{A}_{t}\mathbf{S}_{t}\mathbf{A}_{t}^{\top}\big{\|}_{\mathrm{F}}^{2}-\frac{2}{L\eta}\big{\langle}\mathbf{I}_{m},\mathbf{A}_{t}\mathbf{S}_{t}\mathbf{A}_{t}^{\top}\big{\rangle}_{\mathrm{F}}+\big{\|}\mathbf{B}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\big{\|}_{\mathrm{F}}^{2}-$
@@
=$\overset{(b)}{=}\frac{L\eta^{2}}{2}\|\nabla\ell(\mathbf{W}_{t})\|_{2}^{2}\Big{[}\big{(}\big{\|}\mathbf{A}_{t}\mathbf{S}_{t}\mathbf{A}_{t}^{\top}\big{\|}_{\mathrm{F}}+\big{\|}\mathbf{B}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\big{\|}_{\mathrm{F}}\big{)}^{2}-\frac{2}{L\eta}\times$
@@
=$\overset{(d)}{=}\frac{L\eta^{2}}{2}\|\nabla\ell(\mathbf{W}_{t})\|_{2}^{2}\Big{(}\big{\|}\mathbf{A}_{t}\mathbf{S}_{t}^{\frac{1}{2}}\big{\|}_{\mathrm{F}}^{2}+\big{\|}\mathbf{B}_{t}\mathbf{S}_{t}^{-\frac{1}{2}}\big{\|}_{\mathrm{F}}^{2}-\frac{1}{L\eta}\Big{)}^{2}+\mathcal{O}(L\eta^{3})+\operatorname{Const.}$
@@ B.3 Proof of Theorem 3
Since $f(\mathbf{Q}_{t}^{S},\boldsymbol{\lambda}_{t}^{S})$ is continuous w.r.t. both $\mathbf{Q}_{t}^{S}\in\mathrm{O}(r)$ and $\boldsymbol{\lambda}_{t}^{S}\succeq 0$, its infimum can be determined by checking the limit of $f$ as it approaches the boundary of the open set $\mathbb{S}_{++}^{r}$, and analyzing stationary points in the interior [45].
Notice that $\mathrm{O}(r)$ is closed [45], while $(0,+\infty)^{r}$ is open. It follows from (B.3) that, for any given $\mathbf{Q}_{t}^{S}\in\mathrm{O}(r)$, if some $[\boldsymbol{\lambda}_{t}^{S}]_{i}$ approaches 00 or +$\infty$+ ∞, the objective $f(\mathbf{Q}_{t}^{S},\boldsymbol{\lambda}_{t}^{S})$ goes to +$\infty$+ ∞. As a consequence, the minimum must be attained inside the interior of $\mathbb{S}_{++}^{r}$.
@@
[…]
@@
Given that $g(\mathbf{S}_{t})$ approaches +$\infty$+ ∞ on the boundary of $\mathbb{S}_{++}^{r}$ and has only one stationary point $\tilde{\mathbf{S}}_{t}$, its minimum is reached uniquely at $\tilde{\mathbf{S}}_{t}$, i.e.,
@@
=$\overset{(a)}{=}\operatorname{tr}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t}\tilde{\mathbf{S}}_{t})+\operatorname{tr}(\tilde{\mathbf{S}}_{t}^{-1}\mathbf{B}_{t}^{\top}\mathbf{B}_{t})\overset{(b)}{=}2\operatorname{tr}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t}\tilde{\mathbf{S}}_{t})$
@@
=$\overset{(a)}{=}\operatorname{tr}\big{(}\big{[}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\mathbf{B}_{t}^{\top}\mathbf{B}_{t}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\big{]}^{\frac{1}{2}}\big{)}=\sum_{i=1}^{r}\lambda_{i}\big{(}\big{[}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\mathbf{B}_{t}^{\top}\mathbf{B}_{t}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\big{]}^{\frac{1}{2}}\big{)}$
@@ B.4 Proof of Theorem 5
Plugging $\mathbf{S}_{t}=s_{t}\mathbf{I}_{r},\;s_{t}\in_{++}$ into (8) incurs alternative objective
$\min_{s_{t}\in_{++}}h(s_{t}):=\Big{(}\|\mathbf{A}_{t}\|_{\mathrm{F}}^{2}s_{t}+\|\mathbf{B}_{t}\|_{\mathrm{F}}^{2}s_{t}^{-1}-\frac{1}{L\eta}\Big{)}^{2}.$ (21)
@@
Since $(\cdot)^{2}$ monotonically increases on +, the unique global optima of (21) is
$s_{t}^{*}=\operatorname*{arg\,min}_{s_{t}\in_{++}}\|\mathbf{A}_{t}\|_{\mathrm{F}}^{2}s_{t}+\|\mathbf{B}_{t}\|_{\mathrm{F}}^{2}s_{t}^{-1}-\frac{1}{L\eta}=\frac{\|\mathbf{B}_{t}\|_{\mathrm{F}}}{\|\mathbf{A}_{t}\|_{\mathrm{F}}}.$
@@
where the quartic polynomial in the parenthesis has coefficients with signs (+,,+,)(+,-,+,-)( + , - , + , - ). Using Theorem 8, the gradient (B.4) has $3$ or $1$ positive roots. That says, (21) has either $3$ or $1$ stationary point(s).
Notice that the objective (21) must be non-negative, and its lower bound of 00 can be reached if and only if $\|\mathbf{A}_{t}\|_{\mathrm{F}}^{2}s_{t}+\|\mathbf{B}_{t}\|_{\mathrm{F}}^{2}s_{t}^{-1}-1/(L\eta)=0$. Solving this quadratic equation over $s_{t}\in_{++}$ leads to two global minimum
@@
Lastly, the objective (21) is a continuous function of $s_{t}\in(0,+\infty)$, and tends to +$\infty$+ ∞ when $s_{t}\to 0$ and $s_{t}\to+\infty$. We conclude it only has two global minimum given by (23). ∎

### v2

@@ B.1 Proof of Lemma 1
This suggests, there must exist an invertible matrix $\mathbf{P}_{t}\in\mathbb{R}^{r\times r}$ such that $\tilde{\mathbf{A}}_{t}=\mathbf{A}_{t}\mathbf{P}_{t}$.
@@
$\overset{(a)}{=}-\eta\tilde{\mathbf{A}}_{t}\tilde{\mathbf{A}}_{t}^{\top}\nabla\ell(\mathbf{W}_{t})-\eta\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t}\tilde{\mathbf{B}}_{t}^{\top}+\eta^{2}\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t}\tilde{\mathbf{A}}_{t}^{\top}\nabla\ell(\mathbf{W}_{t})$
$\overset{(b)}{=}-\eta\mathbf{A}_{t}\mathbf{P}_{t}\mathbf{P}_{t}^{\top}\mathbf{A}_{t}^{\top}\nabla\ell(\mathbf{W}_{t})-\eta\nabla\ell(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{P}_{t}^{-\top}\mathbf{P}_{t}^{-1}\mathbf{B}_{t}^{\top}+\eta^{2}\nabla\ell(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{A}_{t}^{\top}\nabla\ell(\mathbf{W}_{t})$
$\overset{(c)}{=}\mathbf{A}_{t}(\mathbf{P}_{t}\mathbf{P}_{t}^{\top})\Delta\mathbf{B}_{t}^{\top}+\Delta\mathbf{A}_{t}(\mathbf{P}_{t}\mathbf{P}_{t}^{\top})^{-1}\mathbf{B}_{t}^{\top}+\Delta\mathbf{A}_{t}\Delta\mathbf{B}_{t}^{\top}$ (12)
where $(a)$ uses (3) and that $\mathbf{W}_{t}^{\mathrm{pt}}+\tilde{\mathbf{A}}_{t}\tilde{\mathbf{B}}_{t}^{\top}=\mathbf{W}_{t}^{\mathrm{pt}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}=\mathbf{W}_{t}$, $(b)$ is due to $\tilde{\mathbf{A}}_{t}=\mathbf{A}_{t}\mathbf{P}_{t}$ and $\tilde{\mathbf{B}}_{t}=\mathbf{B}_{t}\mathbf{P}^{-\top}$, and $(c)$ leverages (1).
@@ B.2 Proof of Proposition 2
$\overset{(a)}{=}\ell(\mathbf{W}_{t})+\big\langle\nabla\ell(\mathbf{W}_{t}),\mathbf{A}_{t}\mathbf{S}_{t}\Delta\mathbf{B}_{t}^{\top}+\Delta\mathbf{A}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}-\Delta\mathbf{A}_{t}\Delta\mathbf{B}_{t}^{\top}\big\rangle_{\mathrm{F}}+$
@@
$\overset{(b)}{=}\big\langle\nabla\ell(\mathbf{W}_{t}),\mathbf{A}_{t}\mathbf{S}_{t}\Delta\mathbf{B}_{t}^{\top}+\Delta\mathbf{A}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\big\rangle_{\mathrm{F}}+\frac{L}{2}\Big[\big\|\mathbf{A}_{t}\mathbf{S}_{t}\Delta\mathbf{B}_{t}^{\top}\big\|_{\mathrm{F}}^{2}+\big\|\Delta\mathbf{A}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\big\|_{\mathrm{F}}^{2}+$
@@
$\overset{(c)}{=}\frac{L}{2}\Big[\Big\|\frac{1}{L}\nabla\ell(\mathbf{W}_{t})+\mathbf{A}_{t}\mathbf{S}_{t}\Delta\mathbf{B}_{t}^{\top}\Big\|_{\mathrm{F}}^{2}+\Big\|\frac{1}{L}\nabla\ell(\mathbf{W}_{t})+\Delta\mathbf{A}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\Big\|_{\mathrm{F}}^{2}+$ (13)
@@
When performing fine-tuning from a pretrained weight, both $\eta$ and $\|\nabla\ell(\mathbf{W}_{t})\|_{2}$ are observed to be tiny [^1] Typically, $\eta=\mathcal{O}(10^{-4})$, and $\|\nabla\ell(\mathbf{W}_{t})\|_{2}\leq\|\nabla\ell(\mathbf{W}_{t})\|_{\mathrm{F}}=\mathcal{O}(10^{-1})$ per matrix.. As a result, (B.2) is dominated by (B.2), and is neglectable in practice; see also LoRA-GA ; LoRA-RITE .
@@
$\overset{(a)}{=}\frac{L\eta^{2}}{2}\|\nabla\ell(\mathbf{W}_{t})\|_{2}^{2}\Big[\big\|\mathbf{A}_{t}\mathbf{S}_{t}\mathbf{A}_{t}^{\top}\big\|_{\mathrm{F}}^{2}-\frac{2}{L\eta}\big\langle\mathbf{I}_{m},\mathbf{A}_{t}\mathbf{S}_{t}\mathbf{A}_{t}^{\top}\big\rangle_{\mathrm{F}}+\big\|\mathbf{B}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\big\|_{\mathrm{F}}^{2}$-
@@
$\overset{(b)}{=}\frac{L\eta^{2}}{2}\|\nabla\ell(\mathbf{W}_{t})\|_{2}^{2}\Big[\big(\big\|\mathbf{A}_{t}\mathbf{S}_{t}\mathbf{A}_{t}^{\top}\big\|_{\mathrm{F}}+\big\|\mathbf{B}_{t}\mathbf{S}_{t}^{-1}\mathbf{B}_{t}^{\top}\big\|_{\mathrm{F}}\big)^{2}-\frac{2}{L\eta}\times$
@@
$\overset{(d)}{=}\frac{L\eta^{2}}{2}\|\nabla\ell(\mathbf{W}_{t})\|_{2}^{2}\Big(\big\|\mathbf{A}_{t}\mathbf{S}_{t}^{\frac{1}{2}}\big\|_{\mathrm{F}}^{2}+\big\|\mathbf{B}_{t}\mathbf{S}_{t}^{-\frac{1}{2}}\big\|_{\mathrm{F}}^{2}-\frac{1}{L\eta}\Big)^{2}+\mathcal{O}(L\eta^{3})+\operatorname{Const.}$
@@ B.3 Proof of Theorem 3
Since $f(\mathbf{Q}_{t}^{S},\boldsymbol{\lambda}_{t}^{S})$ is continuous w.r.t. both $\mathbf{Q}_{t}^{S}\in\mathrm{O}(r)$ and $\boldsymbol{\lambda}_{t}^{S}\succeq 0$, its infimum can be determined by checking the limit of $f$ as it approaches the boundary of the open set $\mathbb{S}_{++}^{r}$, and analyzing stationary points in the interior MathAnalysis-Rudin .
Notice that $\mathrm{O}(r)$ is closed MathAnalysis-Rudin , while $(0,+\infty)^{r}$ is open. It follows from (B.3) that, for any given $\mathbf{Q}_{t}^{S}\in\mathrm{O}(r)$, if some $\boldsymbol{\lambda}_{t}^{S}]_{i}$ approaches 0 or $\infty$, the objective $f(\mathbf{Q}_{t}^{S},\boldsymbol{\lambda}_{t}^{S})$ goes to $\infty$. As a consequence, the minimum must be attained inside the interior of $\mathbb{S}_{++}^{r}$.
@@
$\Rightarrow$~~ $\mathbf{S}_{t}\mathbf{A}_{t}^{\top}\mathbf{A}_{t}\mathbf{S}_{t}=\mathbf{B}_{t}^{\top}\mathbf{B}_{t}$ (19)
$\Rightarrow$~~ $(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\mathbf{S}_{t}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\mathbf{S}_{t}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}=(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\mathbf{B}_{t}^{\top}\mathbf{B}_{t}^{\top}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}$
$\Rightarrow$~~ $\big[(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\mathbf{S}_{t}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\big]^{2}=(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\mathbf{B}_{t}^{\top}\mathbf{B}_{t}^{\top}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}.$
@@
Given that $g(\mathbf{S}_{t})$ approaches $\infty$ on the boundary of $\mathbb{S}_{++}^{r}$ and has only one stationary point $\tilde{\mathbf{S}}_{t}$, its minimum is reached uniquely at $\tilde{\mathbf{S}}_{t}$, i.e.,
@@
$\overset{(a)}{=}\operatorname{tr}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t}\tilde{\mathbf{S}}_{t})+\operatorname{tr}(\tilde{\mathbf{S}}_{t}^{-1}\mathbf{B}_{t}^{\top}\mathbf{B}_{t})\overset{(b)}{=}2\operatorname{tr}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t}\tilde{\mathbf{S}}_{t})$
@@
$\overset{(a)}{=}\operatorname{tr}\big(\big[(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\mathbf{B}_{t}^{\top}\mathbf{B}_{t}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\big]^{\frac{1}{2}}\big)=\sum_{i=1}^{r}\lambda_{i}\big(\big[(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\mathbf{B}_{t}^{\top}\mathbf{B}_{t}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{\frac{1}{2}}\big]^{\frac{1}{2}}\big)$
@@ B.4 Proof of Theorem 5
Plugging $\mathbf{S}_{t}=s_{t}\mathbf{I}_{r},\;s_{t}\in\mathbb{R}_{++}$ into (8) incurs alternative objective
$\min_{s_{t}\in\mathbb{R}_{++}}h(s_{t}):=\Big(\|\mathbf{A}_{t}\|_{\mathrm{F}}^{2}s_{t}+\|\mathbf{B}_{t}\|_{\mathrm{F}}^{2}s_{t}^{-1}-\frac{1}{L\eta}\Big)^{2}.$ (21)
@@
Since $(\cdot)^{2}$ monotonically increases on $\mathbb{R}_{+}$, the unique global optimum of (21) is
$s_{t}^{*}=\operatorname*{arg\,min}_{s_{t}\in\mathbb{R}_{++}}\|\mathbf{A}_{t}\|_{\mathrm{F}}^{2}s_{t}+\|\mathbf{B}_{t}\|_{\mathrm{F}}^{2}s_{t}^{-1}-\frac{1}{L\eta}=\frac{\|\mathbf{B}_{t}\|_{\mathrm{F}}}{\|\mathbf{A}_{t}\|_{\mathrm{F}}}.$
@@
where the quartic polynomial in the parenthesis has coefficients with signs (+,,+,)(+,-,+,-). Using Theorem 8, the gradient (B.4) has $3$ or $1$ positive roots. That says, (21) has either $3$ or $1$ stationary point(s).
Notice that the objective (21) must be non-negative, and its lower bound of 0 can be reached if and only if $\|\mathbf{A}_{t}\|_{\mathrm{F}}^{2}s_{t}+\|\mathbf{B}_{t}\|_{\mathrm{F}}^{2}s_{t}^{-1}-1/(L\eta)=0$. Solving this quadratic equation over $s_{t}\in\mathbb{R}_{++}$ leads to two global minimum
@@
Lastly, the objective (21) is a continuous function of $s_{t}\in(0,+\infty)$, and tends to $\infty$ when $s_{t}\to 0$ and $s_{t}\to+\infty$. We conclude it only has two global minimum given by (23). ∎